What is the pH of the solution created when 1.00 mL of NaOH is titrated into 10 mL of 0.1 M acetic acid and acetate buffer solution?

Concentration of NaOH is 0.220849 M.
Answer #1

no of moles of CH3COOH   = molarity * volume in L

                                         = 0.1*0.01   = 0.001moles

no of moles of NaOH    = molarity * volume in L

                                           = 0.220849*0.001    = 0.000220849moles

PKa = 4.75

------------ CH3COOH(aq)    +    NaOH(aq) -------------> CH3COONa(aq) + H2O(l)

I------------- 0.001---------------------- 0.000220849 ------------------ 0

C----------- -0.000220849 ---------- -0.000220849-------------------0.000220849

E------------ 0.000779151-------------- 0 -------------------------------0.000220849

     PH   = PKa + log[CH3COONa]/[CH3COOH]

             = 4.75 + log0.000220849/0.000779151

               = 4.75 -0.5475

               = 4.2025 >>>>>answer